KTH Matematik Lars Filipsson

Grupparbete 4 i kursen Amelia 1 för P och T ht 04

Inlämning 12/11. Ger maximalt 6 momentpoäng.

{y=x + 2x = 1_{\overline{x^{2}−4x+5}} \atop =xln 4x.}

Is this < a, z > or ⟨a, z⟩.